PLEASE ANSWER QUICKLY!

1.True or False - Each stock's rate of return in a given year consists of a dividend yield (which might be zero) plus a capital gains yield (which could be positive, negative, or zero). Such returns are calculated for all the stocks in the S&P 500. A simple average of those returns (which gives equal weight to each company in the S&P 500) is then calculated. That average is called the “return on the S&P Index,” and it is often used as an indicator of the “return on the market.”

2. If you save $3,000 a year, and earn 6.5% interest annually, how much will you have in your account after 25 years?

3. What is the present value today of $500 received 6 years from now if the discount rate is 6% compounded daily?

4. The less frequent the compounding the

A.

greater the amount deposited.

B.

greater the future value.

C.

lower the effective interest rate.

D.

lower the present value.

5. Sarah bought a raft for $608.66 using her credit card. The interest rate is 18% compounded monthly. If she pays $10 a month, how long will it take her to pay off the credit card?

6. Find the Effective Interest Rate for the Stated Rate of 0.05 compounded 9 times per year.

7. True of False - If an investor in a common stock expects a return of 6% a year on his investment, he is expecting a dividend yield of 6% a year.

8. Star Products Company has purchased a piece of equipment costing $200,000. It is expected to generate a ten year stream of cash flows amounting to $23,273 per year. Determine the rate of return they expect to earn from this purchase.

9. True of False - If a company uses a stock split, the owners of common stock do not lose money because the stock is now priced less.
Answer #1

As per rules I am answering the first 4 subparts of the question

1: True

S&P Return includes dividend and capital gains.

2: 36593.63

PV of annuity = Annuity*(1-1/(1+rate)^number of terms)/rate

= 3000*(1-1/(1+6.5%)^25)/6.5%

=36593.63

3: 348.85

Present value =Future value/(1+Rate)^number of periods

= 500/(1+6%/365)^(6*365)

=348.85

4: C

Due to lesser compounding the effective rate will be lower

EAR= (1+APR/m)^m-1 where m is the number of compounding periods.